a car leaves town and drives east at 50 miles per hour. a second car leaves the same point three hours later and also drives north at 50 miles per hour. how far apart are they 4 hours after the second care leaves?
Answer by richwmiller(17219)

The first car went 3*50=150 miles before the second car left
after 4 more hours they each go 200 miles.
So the first car has gone 350 and the second car 200
the distance between them d=sqrt(350^2+200^2)=50sqrt(65)=403.11